OpenStudy (anonymous):

hi guys can you solve this using inverse laplace "2 over (s+1)^2 times (s^2 +4)

OpenStudy (ash2326):

\[\frac{2}{(s+1)^2(s^2+4)}\] We need to find the partial fractions for this. Do you know how to do that?

OpenStudy (anonymous):

i just did but i cant find the abcd help me pls

OpenStudy (ash2326):

I never use the A, B, C method. Let me show you how I do this.

OpenStudy (anonymous):

teach me master

OpenStudy (ash2326):

Subtract the two terms in the denominator, \[(s^2+4)-(s+1)^2=\underline{-2s+3}\] I'll create the underlined term in the numerator \[\frac{2}{3}\frac{3}{(s+1)^2(s^2+4)}\] \[\frac{2}{3}(\frac{3-2s+2s}{(s+1)^2(s^2+4)})\] Do you get this part?

OpenStudy (anonymous):

where did you get 2/3

OpenStudy (ash2326):

I had a 2 in the numerator, I took it out then I multiplied and divided by 3 \[\frac{2}{1}\] \[\frac{2\times 3}{3}=\frac{2}{3} \times 3\]

OpenStudy (anonymous):

ahh ok get it just continue to answer ur own step

OpenStudy (ash2326):

Cool :)

OpenStudy (ash2326):

Wait a minute, the steps I took, aren't helping. Just a moment. I'll think for sometime

OpenStudy (anonymous):

awts

OpenStudy (ash2326):

OK, Let's use the partial fraction method here. My method is failing here \[\frac 2 {(s+1)^2(s^2+4)}=\frac A{(s+1)}+\frac B {(s+1)^2}+\frac{Cs+D}{s^2+4}\]

OpenStudy (ash2326):

Let's rearrange this, we'd get \[2=A(s+1)(s^2+4)+B(s^2+4)+(Cs+D)(s+1)^2\] First put s=-1 \[2=B(1^2+4)\]or \[B=\frac 2 5\] Do you get it @coldi3845 ?

OpenStudy (anonymous):

but 2Cs right? bcoz the derivative of s^2 +4 is 2S?

OpenStudy (ash2326):

We can do that way also, here C will come twice of what we get if we use (2Cs)

OpenStudy (anonymous):

how did you get=-1

OpenStudy (ash2326):

I didn't get, I put s=-1, which makes the first and the last term on right side 0, then I can find the value of B

OpenStudy (ash2326):

@coldi3845 do you get this?

OpenStudy (anonymous):

not really

OpenStudy (ash2326):

OK, I'll explain you with this \[\frac 1{(s+1)(s+2)}=\frac A{s+1}+\frac B{s+2}\] We need to find the value of A and B.

OpenStudy (ash2326):

@coldi3845 ???

OpenStudy (anonymous):

why it becomes A over s+1 and B over s+2

OpenStudy (ash2326):

We are splitting the fraction on the right to 2 partial fractions. It'd be easy to fine the inverse laplace transform that way. isn't it?

OpenStudy (anonymous):

ok ok just continue

OpenStudy (ash2326):

Should I continue, with our question? Do you get how I find the value of B?

OpenStudy (anonymous):

yeah u assume s=-1 to find it

OpenStudy (ash2326):

\[2=A(s+1)(s^2+4)+B(s^2+4)+(Cs+D)(s+1)^2\] now Let's put s=+2i that will make \[s^2+4=-4+4=0\] \[2=(C\times 2i+D)(-4+1+4i)\] \[2=(C\times 2i+D)(-3+4i)\] \[2=-6iC-8C-3D+4iD\] Equate imaginary and real parts \[-8C-3D=2\] \[-6C+4D=0\] Do you get this?

OpenStudy (ash2326):

@coldi3845 ???

OpenStudy (anonymous):

ok ok and then?

OpenStudy (ash2326):

Can you find the value of C and D? \[-8C-3D=2\] \[-6C+4D=0\]

OpenStudy (ash2326):

@coldi3845

OpenStudy (anonymous):

no

OpenStudy (ash2326):

You have two equations, two variables. It's simultaneous linear equation @coldi3845

OpenStudy (anonymous):

just finish it sir bcoz i will compare it to my convolution if it is right

OpenStudy (ash2326):

@coldi3845 I can't you the full solution, just like that. I want your active involvement in the solution

OpenStudy (ash2326):

*give

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

just solved it pls